OpenStudy (teller):

Is - SQRT 22 Irrotational or Rotational ?

OpenStudy (teller):



OpenStudy (dinamix):

irrotational i think

OpenStudy (zzr0ck3r):

\(-\sqrt{22}=-\sqrt{11*2}=-\sqrt{2}\sqrt{11}\) Now, the square root of a prime number is irrational, and two irrational numbers multiplied together gives an irrational number.

OpenStudy (teller):

Thank you !